OpenStudy (anonymous):

Which number would you add to 4 to give an approximate sum of 5? 3--- 10 2 2-- 3 ------------ 6 1-- 11 ------------ 1 1-- 12 ------------ 7 2-- 8

OpenStudy (anonymous):

Number one right?!?!

OpenStudy (anonymous):

3 4/10 = 34/10 34/10 + x = 5 x = 50-34/10 = 17/10 = 1 7/10 closest to B 1 6/11